A 10-year loan in the amount of $238,000 is to be repaid in equal annual payments. The interest rate is 7 percent, compounded annually. What is the amount of interest that is included in the loan payment for Year 3? PLEASE HELP WITH EXCEL FUNCTION
Answer #1

C7 - fx fx =PMT(C2,C3,-C4,C5,C6) PMI D B F 0.07 A 1 2 RATE 3 NPER 4 PV 5 FV 6 TYPE 7 PMT 238000 Rate Per Period Total No.of P

C7 fx =PMT(C2,C3,-C4,C5,C6) B А 10.07 10 [238000 Rate Per Period Total No.of Payments Present Value Future value Ending(0), B